OpenStudy (anonymous):

Experimental or Theoretical Probability? You toss a coin 50 times and you end up with 20 heads and 30 tails P(H) = 20/50 = 0.4 P(T) = 30/50 = 0.6 Experimental Theoretical Cannot be determined Neither

OpenStudy (campbell_st):

you have stated the relative frequency of experimental probability. experimental probability is what actually happened when you tossed a coin. Theoretical probability is what should happen P(H) = 25/50 or 1/2 P(T) = 25/50 or 1/2 hope it helps

OpenStudy (anonymous):

so wats the answer